# BasicAVSR: Arbitrary-Scale Video Super-Resolution via Image Priors and Enhanced Motion Compensation

### v1

@@ 3.3 Second-Order Motion Compensation Unit
Our alignment is inherently a two-step refinement process. We first perform the ‘coarse estimation’ using a pre-trained optical flow network to predict the initial optical flows $\bm{f}$ between the current state $\bm{h}_{c}$ and the neighbouring state $\bm{h}_{n}$. This initial optical flow estimation roughly determines the spatial displacement range and establishes a local region of interest for the next step. Following the ‘coarse estimation’, we perform a ‘fine-grained search-based matching’ to obtain the precise sub-pixel offset. To help the network better capture spatial information and enhance feature discrimination, we utilize a position encoding network, which takes coordinates as inputs to model signals. This network serves as a prior for the ‘fine estimation’, with the prior encoded as trainable weights within a MLP. MLPs are theoretically universal approximators capable of representing any function and frequency [20]. Especially showing strong learning ability for high-frequency content [41]. This matching process is formulated as an attention mechanism. Given the coordinate $p$, the aligned feature $\bm{h}_{n\rightarrow c}$ at the spatial position $p$ is aggregated by computing the similarity between the query patch computed by $\bm{h}_{c}$ at $p$ and the key patches computed by $\bm{h}_{n}$ in neighbors $p^{\prime}$. The computation for ‘fine-grained search-based matching’ is as follows:
@@
To further enhance the fidelity of the feature propagation, we adopt a second-order deformable alignment strategy inspired by BasicVSR++ [7]. This approach aligns adjacent two frames with the current frame, introducing a residual deformable convolution. This convolution learns residual offsets $\bm{o}$ and modulation masks $\bm{m}$ to locally correct the minor misalignments that previous steps might have missed, ensuring the final feature aggregation is accurately aligned. Finally, we use the $i$-th frame ($i\textgreater 1$) as an example to demonstrate second-order motion compensation, replacing Eq. (2) with the following omputation formula:
@@ 3.4 Hyper-Upsampling Unit
Inspired by the neural kriging upsampler [59], our hyper-upsampling unit consists of two branches: SR feature preparation and SR kernel prediction, as shown in Fig. 2. For SR feature preparation, we pass the final aligned hidden state $\bm{g}_{i}$ through a ResNet with $N$ residual blocks to compute SR features. Next, we unfold a $K\times K$ spatial neighborhood of $C$-dimensional SR feature representations into $C\times K^{2}$ channels (i.e., the tensor generalization of $\mathtt{img2col}(\cdot)$ in image processing). Finally, we upsample the unfolded features to the target resolution using bilinear interpolation, resulting in $\bm{s}_{i}$.
For SR kernel generation, we train a hyper-network, i.e., a multi-layer perceptron (MLP) with periodic activation functions [9], to predict the upsampling kernels $\bm{w}$. Periodic activations have been shown to effectively address the spectral bias of MLPs, outperforming ReLU non-linearity [52]. The inputs to the MLP are carefully selected to be scale-aware and content-independent. These include 1) the scaling factors $(\alpha,\beta)$, 2) the relative coordinates between the LR and HR frames $(\delta_{\alpha},\delta_{\beta})$, and 3) the spatial indices $(k_{1},k_{2})$ of $\bm{w}$. The first two inputs have been used in other continuous representation methods [10, 30]. To enhance the discriminability of scale-relevant inputs, we employ sinusoidal positional encoding as a pre-processing step. It is noteworthy that our upsampling kernels $\bm{w}$ can be pre-computed and stored for various target resolutions, which accelerates inference time.
@@ 3.5 Architectural Variants: Adapting to Diverse Application Scenarios
In diverse VSR scenarios, bidirectional RNNs require processing and storing bidirectional hidden states for entire video sequences, which demands significant memory and makes such architectures unsuitable for online applications. To meet online requirements, we extend the BasicAVSR framework to a unidirectional RNN variant by removing the backward RNN and keeping only the forward RNN. For scenarios allowing slight latency, we integrate a lookahead strategy that leverages limited future frames, drawing on the strategy used in our previous work ST-AVSR [49]. Specifically, in the ST-AVSR, the alignment strategy and priors are replaced with the proposed image priors and enhanced motion compensation, respectively. In a unidirectional RNN, only the hidden state of the previous frame is stored, and it is overwritten by the new hidden state after computing the current frame, enabling online output. For the unidirectional RNN with lookahead, in addition to the hidden state of the previous frame, it also requires the hidden states of the next $L$ frames. Thus, the total number of stored hidden states is $L+1$, resulting in an output delay of $L$ frames. A bidirectional RNN needs to store the bidirectional hidden states of the entire video sequence, with the total number of stored hidden states being $2T$, making it only suitable for offline processing. All variants achieve state-of-the-art performance (please see Sec 4.3 for detailed analysis).
@@
EDVR [58] 36.03/0.961/0.072 32.59/0.904/0.108 30.24/0.853/0.202 27.02/0.733/0.349 25.38/0.678/0.411
ArbSR [57] 34.48/0.942/0.096 30.51/0.862/0.200 28.38/0.799/0.295 26.32/0.710/0.428 25.08/0.641/0.492
EQSR [59] 34.71/0.943/0.082 30.71/0.867/0.194 28.75/0.804/0.283 26.53/0.718/0.391 25.23/0.645/0.459
RDN [61] LTE [30] 34.63/0.942/0.093 30.64/0.865/0.204 28.65/0.801/0.289 26.46/0.714/0.410 25.15/0.660/0.488
CLIT [9] 34.63/0.942/0.092 30.63/0.865/0.204 28.63/0.801/0.290 26.43/0.714/0.400 25.14/0.661/0.467
OPE [53] 34.05/0.939/0.082 30.52/0.864/0.199 28.63/0.800/0.293 26.37/0.711/0.421 25.04/0.655/0.504
GaussianSR [21] 34.25/0.940/0.091 30.56/0.866/0.201 28.64/0.800/0.291 26.40/0.712/0.419 25.08/0.657/0.501
ContinuousSR [45] —/—/— 30.65/0.866/0.198 28.67/0.801/0.289 26.49/0.715/0.402 25.14/0.662/0.470
SwinIR [32] LTE [30] 34.73/0.943/0.091 30.73/0.866/0.200 28.75/0.804/0.284 26.56/0.718/0.403 25.24/0.669/0.480
CLIT [9] 34.63/0.942/0.093 30.64/0.865/0.205 28.64/0.802/0.291 26.45/0.715/0.400 25.15/0.662/0.466
OPE [53] 33.39/0.935/0.081 29.40/0.820/0.217 28.49/0.785/0.292 26.30/0.698/0.398 25.01/0.648/0.487
GaussianSR [21] 34.31/0.941/0.089 30.60/0.867/0.199 28.69/0.802/0.290 26.42/0.713/0.416 25.08/0.659/0.498
ContinuousSR [45] —/—/— 30.75/0.868/0.197 28.68/0.805/0.287 26.58/0.720/0.401 25.26/0.670/0.467
VideoINR [12] 31.59/0.900/0.144 30.04/0.852/0.197 28.13/0.791/0.263 25.27/0.687/0.374 23.46/0.619/0.470
MoTIF [11] 31.03/0.898/0.100 30.44/0.862/0.186 28.77/0.807/0.260 25.63/0.698/0.369 25.12/0.664/0.467
BF-STVSR [25] 32.06/0.908/0.092 31.38/0.877/0.146 29.29/0.837/0.200 25.98/0.718/0.321 25.42/0.670/0.459
SAVSR [31] 35.66/0.955/0.046 32.19/0.918/0.100 30.61/0.872/0.138 27.03/0.791/0.250 25.59/0.716/0.312
ST-AVSR [49] 36.91/0.969/0.041 33.41/0.937/0.066 31.03/0.897/0.114 27.89/0.812/0.222 26.04/0.746/0.298
@@
ArbSR [57] 25.86/0.815/0.224 24.01/0.721/0.313 21.23/0.540/0.478 20.34/0.515/0.498
EQSR [59] 26.24/0.826/0.210 24.16/0.730/0.300 21.72/0.573/0.443 20.81/0.528/0.472
RDN [61] LTE [30] 25.98/0.818/0.226 24.03/0.722/0.312 21.64/0.565/0.455 20.60/0.522/0.480
CLIT [9] 25.83/0.815/0.223 23.94/0.721/0.312 21.62/0.563/0.458 20.57/0.520/0.491
OPE [53] 25.77/0.818/0.217 23.98/0.719/0.317 21.60/0.559/0.483 20.55/0.528/0.495
GaussianSR [21] 25.81/0.817/0.222 23.99/0.720/0.313 21.61/0.560/0.460 20.56/0.520/0.484
ContinuousSR [45] 25.94/0.820/0.216 24.08/0.725/0.310 21.66/0.568/0.453 20.69/0.525/0.473
SwinIR [32] LTE [30] 26.43/0.826/0.217 24.09/0.727/0.305 21.72/0.570/0.448 20.70/0.524/0.475
CLIT [9] 25.89/0.818/0.224 24.00/0.724/0.314 21.65/0.565/0.457 20.69/0.522/0.479
OPE [53] 25.55/0.801/0.221 23.93/0.711/0.320 21.58/0.521/0.471 20.65/0.520/0.492
GaussianSR [21] 25.92/0.820/0.220 24.01/0.722/0.311 21.63/0.563/0.455 20.66/0.523/0.480
ContinuousSR [45] 26.54/0.830/0.210 24.16/0.729/0.301 21.76/0.573/0.444 20.80/0.539/0.469
VideoINR [12] 23.02/0.715/0.203 24.34/0.741/0.249 20.80/0.536/0.431 20.43/0.511/0.453
MoTIF [11] 23.55/0.734/0.209 24.52/0.746/0.261 20.94/0.546/0.426 20.48/0.518/0.450
BF-STVSR [25] 24.12/0.745/0.166 24.90/0.784/0.222 21.13/0.579/0.423 20.59/0.537/0.447
SAVSR [31] 27.82/0.875/0.088 25.97/0.835/0.154 21.42/0.645/0.359 20.73/0.588/0.393
ST-AVSR [49] 29.09/0.913/0.069 26.16/0.852/0.127 21.60/0.668/0.306 20.64/0.609/0.357
@@ 4.1.1 Datasets
BasicAVSR is trained on the REDS dataset [42], which comprises $240$ videos of resolution $720\times 1,280$ captured by GoPro. Each video consists of $100$ HR frames. Following the settings in [12, 9, 11], we generate LR frames using the bicubic degradation model, with randomly sampled scaling factors $(\alpha,\beta)$ from a uniform distribution $\mathcal{U}[1,4]$. We test BasicAVSR on the validation set of REDS comprising $30$ videos, and the Vid4 dataset [36] containing $4$ videos. To evaluate the generalization of our method to unseen degradation models, we applied a video random degradation pipeline [8] to the test set of GoPro [43], incorporating noise and video compression to synthesize unseen degradations for validation. Additionally, we also use real-world and online collected data to verify the generalization of our method.
@@ 4.1.2 Data Pre-processing for Training
To enable mini-batch training with varying LR/HR resolutions, we adapt the pre-processing method used for AISR in EQSR [59] to AVSR. Specifically, from an HR video patch of size $H\times W\times T$, we generate the input LR video patch by resizing it to $P\times P\times T$. We next crop a set of ground-truth patches of size $P\times P\times T$ from the same HR patch. The respective relative coordinates $(\delta_{\alpha},\delta_{\beta})$ are recorded for use in the hyper-upsampling unit to differentiate between different ground-truth patches for the same input (see the data pre-processing pipeline in Fig. 4). Data augmentation techniques include random rotation (by $90^{\circ}$, $180^{\circ}$, or $270^{\circ}$) and random horizontal and vertical flipping.
@@ 4.1.3 Implementation Details
BasicAVSR is end-to-end optimized for 300300K iterations. Adam [27] is chosen as the optimizer, with an initial learning rate $2\times$10-4 that is gradually lowered to $\times1$0-6 by cosine annealing [37]. We set the input patch size to $P=80$, the sequence length to $T=15$, the number of ResBlocks to $N=5$, the searching window size to $r=2$, the unfolding neighborhood to $K=3$, and the SR feature dimension to $C=64$, respectively. The hidden dimensions of the MLP in the hyper-upsampling unit are $16$, $16$, $16$, and $64$, respectively. The parameters of SPYNet [47] as the optical flow estimator are frozen during training. We use the Charbonnier loss [29]:

### v2

@@ 3.3 Second-Order Motion Compensation Unit
Our alignment is inherently a two-step refinement process. We first perform the ‘coarse estimation’ using a pre-trained optical flow network to predict the initial optical flows $\bm{f}$ between the current state $\bm{h}_{c}$ and the neighbouring state $\bm{h}_{n}$. This initial optical flow estimation roughly determines the spatial displacement range and establishes a local region of interest for the next step. Following the ‘coarse estimation’, we perform a ‘fine-grained search-based matching’ to obtain the precise sub-pixel offset. To help the network better capture spatial information and enhance feature discrimination, we utilize a position encoding network, which takes coordinates as inputs to model signals. This network serves as a prior for the ‘fine estimation’, with the prior encoded as trainable weights within an MLP. MLPs are theoretically universal approximators capable of representing any function and frequency [hornik1989multilayer]. Especially showing strong learning ability for high-frequency content [mildenhall2021nerf]. This matching process is formulated as an attention mechanism. Given the coordinate $p$, the aligned feature $\bm{h}_{n\rightarrow c}$ at the spatial position $p$ is aggregated by computing the similarity between the query patch computed by $\bm{h}_{c}$ at $p$ and the key patches computed by $\bm{h}_{n}$ in neighbors $p^{\prime}$. The computation for ‘fine-grained search-based matching’ is as follows:
@@
To further enhance the fidelity of the feature propagation, we adopt a second-order deformable alignment strategy inspired by BasicVSR++ [chan2022basicvsr++]. This approach aligns adjacent two frames with the current frame, introducing a residual deformable convolution. This convolution learns residual offsets $\bm{o}$ and modulation masks $\bm{m}$ to locally correct the minor misalignments that previous steps might have missed, ensuring the final feature aggregation is accurately aligned. Finally, we use the $i$-th frame ($i\textgreater 1$) as an example to demonstrate second-order motion compensation, replacing Eq. (2) with the following computation formula:
@@ 3.4 Hyper-Upsampling Unit
Inspired by the neural kriging upsampler [wang2023deep], our hyper-upsampling unit consists of two branches: SR feature preparation and SR kernel prediction, as shown in Fig. 2. For SR feature preparation, we pass the final aligned hidden state $\bm{g}_{i}$ through a ResNet with $N$ residual blocks to compute SR features. Next, we unfold a $K\times K$ spatial neighborhood of $C$-dimensional SR feature representations into $C\times K^{2}$ channels (i.e., the tensor generalization of $\mathtt{img2col}(\cdot)$ in image processing). Finally, we upsample the unfolded features to the target resolution using bilinear interpolation, resulting in $\bm{s}_{i}$.
For SR kernel generation, we train a hyper-network, i.e., a multi-layer perceptron (MLP) with periodic activation functions [chen2023cascaded], to predict the upsampling kernels $\bm{w}$. Periodic activations have been shown to effectively address the spectral bias of MLPs, outperforming ReLU non-linearity [sitzmann2020implicit]. The inputs to the MLP are carefully selected to be scale-aware and content-independent. These include 1) the scaling factors $(\alpha,\beta)$, 2) the relative coordinates between the LR and HR frames $(\delta_{\alpha},\delta_{\beta})$, and 3) the spatial indices $(k_{1},k_{2})$ of $\bm{w}$. The first two inputs have been used in other continuous representation methods [chen2021learning, lee2022local]. To enhance the discriminability of scale-relevant inputs, we employ sinusoidal positional encoding as a pre-processing step. It is noteworthy that our upsampling kernels $\bm{w}$ can be pre-computed and stored for various target resolutions, which accelerates inference time.
@@ 3.5 Architectural Variants: Adapting to Diverse Application Scenarios
In diverse VSR scenarios, bidirectional RNNs require processing and storing bidirectional hidden states for entire video sequences, which demands significant memory and makes such architectures unsuitable for online applications. To meet online requirements, we extend the BasicAVSR framework to a unidirectional RNN variant by removing the backward RNN and keeping only the forward RNN. For scenarios allowing slight latency, we integrate a lookahead strategy that leverages limited future frames, drawing on the strategy used in our previous work ST-AVSR [shang2024arbitrary]. Specifically, in the ST-AVSR, the alignment strategy and priors are replaced with the proposed image priors and enhanced motion compensation, respectively. In a unidirectional RNN, only the hidden state of the previous frame is stored, and it is overwritten by the new hidden state after computing the current frame, enabling online output. For the unidirectional RNN with lookahead, in addition to the hidden state of the previous frame, it also requires the hidden states of the next $L$ frames. Thus, the total number of stored hidden states is $L+1$, resulting in an output delay of $L$ frames. A bidirectional RNN needs to store the bidirectional hidden states of the entire video sequence, with the total number of stored hidden states being $2T$, making it only suitable for offline processing. All variants achieve state-of-the-art performance (please see Sec 4.3 for detailed analysis).
@@
EDVR [wang2019edvr] 36.03/0.961/0.072 32.59/0.904/0.108 30.24/0.853/0.202 27.02/0.733/0.349 25.38/0.678/0.411
ArbSR [wang2021learning] 34.48/0.942/0.096 30.51/0.862/0.200 28.38/0.799/0.295 26.32/0.710/0.428 25.08/0.641/0.492
EQSR [wang2023deep] 34.71/0.943/0.082 30.71/0.867/0.194 28.75/0.804/0.283 26.53/0.718/0.391 25.23/0.645/0.459
RDN [zhang2018residual] LTE [lee2022local] 34.63/0.942/0.093 30.64/0.865/0.204 28.65/0.801/0.289 26.46/0.714/0.410 25.15/0.660/0.488
CLIT [chen2023cascaded] 34.63/0.942/0.092 30.63/0.865/0.204 28.63/0.801/0.290 26.43/0.714/0.400 25.14/0.661/0.467
OPE [song2023ope] 34.05/0.939/0.082 30.52/0.864/0.199 28.63/0.800/0.293 26.37/0.711/0.421 25.04/0.655/0.504
GaussianSR [hu2025gaussiansr] 34.25/0.940/0.091 30.56/0.866/0.201 28.64/0.800/0.291 26.40/0.712/0.419 25.08/0.657/0.501
ContinuousSR [peng2025pixel] —/—/— 30.65/0.866/0.198 28.67/0.801/0.289 26.49/0.715/0.402 25.14/0.662/0.470
SwinIR [liang2021swinir] LTE [lee2022local] 34.73/0.943/0.091 30.73/0.866/0.200 28.75/0.804/0.284 26.56/0.718/0.403 25.24/0.669/0.480
CLIT [chen2023cascaded] 34.63/0.942/0.093 30.64/0.865/0.205 28.64/0.802/0.291 26.45/0.715/0.400 25.15/0.662/0.466
OPE [song2023ope] 33.39/0.935/0.081 29.40/0.820/0.217 28.49/0.785/0.292 26.30/0.698/0.398 25.01/0.648/0.487
GaussianSR [hu2025gaussiansr] 34.31/0.941/0.089 30.60/0.867/0.199 28.69/0.802/0.290 26.42/0.713/0.416 25.08/0.659/0.498
ContinuousSR [peng2025pixel] —/—/— 30.75/0.868/0.197 28.68/0.805/0.287 26.58/0.720/0.401 25.26/0.670/0.467
VideoINR [chen2022videoinr] 31.59/0.900/0.144 30.04/0.852/0.197 28.13/0.791/0.263 25.27/0.687/0.374 23.46/0.619/0.470
MoTIF [chen2023motif] 31.03/0.898/0.100 30.44/0.862/0.186 28.77/0.807/0.260 25.63/0.698/0.369 25.12/0.664/0.467
BF-STVSR [kim2025bf] 32.06/0.908/0.092 31.38/0.877/0.146 29.29/0.837/0.200 25.98/0.718/0.321 25.42/0.670/0.459
SAVSR [li2024savsr] 35.66/0.955/0.046 32.19/0.918/0.100 30.61/0.872/0.138 27.03/0.791/0.250 25.59/0.716/0.312
ST-AVSR [shang2024arbitrary] 36.91/0.969/0.041 33.41/0.937/0.066 31.03/0.897/0.114 27.89/0.812/0.222 26.04/0.746/0.298
@@
ArbSR [wang2021learning] 25.86/0.815/0.224 24.01/0.721/0.313 21.23/0.540/0.478 20.34/0.515/0.498
EQSR [wang2023deep] 26.24/0.826/0.210 24.16/0.730/0.300 21.72/0.573/0.443 20.81/0.528/0.472
RDN [zhang2018residual] LTE [lee2022local] 25.98/0.818/0.226 24.03/0.722/0.312 21.64/0.565/0.455 20.60/0.522/0.480
CLIT [chen2023cascaded] 25.83/0.815/0.223 23.94/0.721/0.312 21.62/0.563/0.458 20.57/0.520/0.491
OPE [song2023ope] 25.77/0.818/0.217 23.98/0.719/0.317 21.60/0.559/0.483 20.55/0.528/0.495
GaussianSR [hu2025gaussiansr] 25.81/0.817/0.222 23.99/0.720/0.313 21.61/0.560/0.460 20.56/0.520/0.484
ContinuousSR [peng2025pixel] 25.94/0.820/0.216 24.08/0.725/0.310 21.66/0.568/0.453 20.69/0.525/0.473
SwinIR [liang2021swinir] LTE [lee2022local] 26.43/0.826/0.217 24.09/0.727/0.305 21.72/0.570/0.448 20.70/0.524/0.475
CLIT [chen2023cascaded] 25.89/0.818/0.224 24.00/0.724/0.314 21.65/0.565/0.457 20.69/0.522/0.479
OPE [song2023ope] 25.55/0.801/0.221 23.93/0.711/0.320 21.58/0.521/0.471 20.65/0.520/0.492
GaussianSR [hu2025gaussiansr] 25.92/0.820/0.220 24.01/0.722/0.311 21.63/0.563/0.455 20.66/0.523/0.480
ContinuousSR [peng2025pixel] 26.54/0.830/0.210 24.16/0.729/0.301 21.76/0.573/0.444 20.80/0.539/0.469
VideoINR [chen2022videoinr] 23.02/0.715/0.203 24.34/0.741/0.249 20.80/0.536/0.431 20.43/0.511/0.453
MoTIF [chen2023motif] 23.55/0.734/0.209 24.52/0.746/0.261 20.94/0.546/0.426 20.48/0.518/0.450
BF-STVSR [kim2025bf] 24.12/0.745/0.166 24.90/0.784/0.222 21.13/0.579/0.423 20.59/0.537/0.447
SAVSR [li2024savsr] 27.82/0.875/0.088 25.97/0.835/0.154 21.42/0.645/0.359 20.73/0.588/0.393
ST-AVSR [shang2024arbitrary] 29.09/0.913/0.069 26.16/0.852/0.127 21.60/0.668/0.306 20.64/0.609/0.357
@@ 4.1.1 Datasets
BasicAVSR is trained on the REDS dataset [nah2019ntire], which comprises $240$ videos of resolution $720\times 1,280$ captured by GoPro. Each video consists of $100$ HR frames. Following the settings in [chen2022videoinr, chen2023cascaded, chen2023motif], we generate LR frames using the bicubic degradation model, with randomly sampled scaling factors $(\alpha,\beta)$ from a uniform distribution $\mathcal{U}[1,4]$. We test BasicAVSR on the validation set of REDS comprising $30$ videos, and the Vid4 dataset [liu2013bayesian] containing $4$ videos. To evaluate the generalization of our method to unseen degradation models, we applied a video random degradation pipeline [chan2022investigating] to the test set of GoPro [Nah_2017_CVPR], incorporating noise and video compression to synthesize unseen degradations for validation. Additionally, we also use real-world and online collected data to verify the generalization of our method.
@@ 4.1.2 Data Pre-processing for Training
To enable mini-batch training with varying LR/HR resolutions, we adapt the pre-processing method used for AISR in EQSR [wang2023deep] to AVSR. Specifically, from an HR video patch of size $H\times W\times T$, we generate the input LR video patch by resizing it to $P\times P\times T$. We next crop a set of ground-truth patches of size $P\times P\times T$ from the same HR patch. The respective relative coordinates $(\delta_{\alpha},\delta_{\beta})$ are recorded for use in the hyper-upsampling unit to differentiate between different ground-truth patches for the same input (see the data pre-processing pipeline in Fig. 4). Data augmentation techniques include random rotation (by $90^{\circ}$, $180^{\circ}$, or $270^{\circ}$) and random horizontal and vertical flipping.
@@ 4.1.3 Implementation Details
BasicAVSR is end-to-end optimized for 300300K iterations. Adam [kingma2014adam] is chosen as the optimizer, with an initial learning rate $2\times$10-4 that is gradually lowered to $\times1$0-6 by cosine annealing [loshchilov2016sgdr]. We set the input patch size to $P=80$, the sequence length to $T=15$, the number of ResBlocks to $N=5$, the searching window size to $r=2$, the unfolding neighborhood to $K=3$, and the SR feature dimension to $C=64$, respectively. The hidden dimensions of the MLP in the hyper-upsampling unit are $16$, $16$, $16$, and $64$, respectively. The parameters of SPYNet [ranjan2017optical] as the optical flow estimator are frozen during training. We use the Charbonnier loss [lai2017deep]: